OpenStudy (anonymous):

Write a word problem describing the time it takes to complete an activity individually and with a friend. For example, if John takes 2 hours to mow his lawn and it takes his sister Maria 4 hours to mow the same lawn, how long would it take John and Maria to mow the lawn together?

OpenStudy (anonymous):

@KamiBug @ardentpikachu

OpenStudy (kamibug):

If it takes Kellie 3 hours to clean the house and June cleans it in 5 hours, how long will it take them to do it together? Lol idk :P You could make up anything ^.^

OpenStudy (anonymous):

I'm just lazy cx

OpenStudy (kamibug):

Heh, that'll get you places! xD Lol jk. Best of luck. :3

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

2. Write a rational equation based upon the word problem you created. The rational equation based upon the scenario above would be . Each fraction represents the amount of the lawn mowed in one hour. The fraction of is John’s portion. Maria’s share is represented by .

OpenStudy (anonymous):

This is where I struggle. @KamiBug HELP!!

OpenStudy (kamibug):

Are there parts missing there?

OpenStudy (anonymous):

yes, \[\frac{ 1 }{ 2 }+\frac{ 1 }{ 4}=\frac{ 1 }{ x }\]

OpenStudy (kamibug):

Okiedokey. Well, in the one I made up it would be 

OpenStudy (anonymous):

Now it wants me to solve.

OpenStudy (kamibug):

You need common denominators. C:You can ignore the denominators now and solve the numerators like an equation! ^.^